Solve for x

1.8 x 10-5 = (0.0125 - 0) (0.025 - 2) = 2
Answer #1

18x105 = (0.0125-x) (0.025:-x) - 1.8X105 2 = 0.0003125 -0.0375x + x² az-0.0375x-1.8 x168 +0.0003125 zo 12? -0.037518x + 0.000

เ” 140 = / 0-032512) VX) x 2 - 55 31 - V 6 (0) 906 - 0 - 2 เS - 30 5D DIS+6 - 0: 0125S 1 -(-0637s)* -0124= 0 0 25.94 2%) 1 2